OpenStudy (anonymous):

For f(x)=5x+4, find f(-1)

OpenStudy (anonymous):

-1 Substitute -1 to the x :)

OpenStudy (anonymous):

insert x=-1 in the f(x)

OpenStudy (anonymous):

iis it like this: 5(-1)+4?

OpenStudy (anonymous):

and solve it ....it is very easy

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

and it gives me -9 right? :)

OpenStudy (anonymous):

nooo

OpenStudy (anonymous):

no? :(

OpenStudy (anonymous):

5*(-1)=-5+4=-1

OpenStudy (anonymous):

i put it on my caculator as 5(-1)+4 and it got me -9

OpenStudy (anonymous):

no loll

OpenStudy (anonymous):

-1 is multiplied by 5and thn add 4 in it

OpenStudy (anonymous):

dang it! i thought i had it, well can you show me how you put it in your caculator? lol

OpenStudy (solomonzelman):

yeah, it's correct that you plug in -1 instead of x, but you don't get -9, lol.

OpenStudy (anonymous):

Honestly i'd recommend you not to rely on your calculator for mathematics to some extent. :) Obviously sometimes its necessary but try to keep it down it helps you think faster :)

OpenStudy (anonymous):

ohhh! wow idk how i got -9 at first, i re did it and i got -1

OpenStudy (anonymous):

haha okay

OpenStudy (solomonzelman):

Good that you got -1 :)